OpenStudy (anonymous):

Simplify 10x - 3x + (-5x).

pooja195 (pooja195):

Same thing combine like terms

pooja195 (pooja195):

10-3-5

OpenStudy (anonymous):

so 2x?

OpenStudy (haseeb96):

yes correct

OpenStudy (anonymous):

or is it nagative?

pooja195 (pooja195):

Yes :D you are getting the hang of it :) no thats right

OpenStudy (haseeb96):

no it will be +ve because greatest value is positive now you have been able to solve these kind of problems

pooja195 (pooja195):

10-5=5-3=2

OpenStudy (anonymous):

thanks